OpenStudy (anonymous):

Can you show me a step by step way to add this fraction? 5/6+3/5. I need to know how to do this problem by using LCM or GCF. Can you do this and show me all the steps?

OpenStudy (swissgirl):

OK so for starter we need to find a common denomanator So our denomanators are 6 and 5 in this case

OpenStudy (swissgirl):

So we must find LCM

OpenStudy (swissgirl):

So lets list the multiples of 5 and 6 5=5,10,15,20,25,30,35 6=6,12,18,24,30 So as we see our LCM is 30

OpenStudy (swissgirl):

Did you follow so far?

OpenStudy (swissgirl):

U didnt respond. U have to type yes in the text and the press post its a blue button on the bottom right hand side of the textbox

OpenStudy (anonymous):

yes I finally see your message.. thanks... this is where I get confused, once I get to this part. Are you still here

OpenStudy (swissgirl):

Yaaaaaa

OpenStudy (swissgirl):

Ok so now we know that our commone denomanator will be 30 correct

OpenStudy (anonymous):

The last thing I seen was the LCM message

OpenStudy (anonymous):

yes.. because they both has 30 as a LCM

OpenStudy (swissgirl):

okkkk

OpenStudy (anonymous):

Great!

OpenStudy (swissgirl):

\( \huge \frac{5} {6} * \frac{5} {5} + \frac {3} {5} *\frac{6} {6} \) \(\huge \frac{25} {30} + \frac{18} {30} =?\)

OpenStudy (anonymous):

Ok, I'm following you, so fa

OpenStudy (anonymous):

far

OpenStudy (swissgirl):

ok so basically in in order to get the denomanater to 30 of the first term I had to multiply it by 5 but whatever i do to the bottom I have to do to the top \(\large \frac{5} {6} * \frac{5} {5} = \frac{25} {30} \) and the second term I had to multiply it by 6 in order to get the denomanator to 30 \( \large \frac{3} {5} * \frac{6} {6} = \frac {18} {30} \) and now all you gotta do is add the 2 fractions So when adding fractions, we keep the denomanator and and add the numerator \( \huge \frac{25} {30}+\frac {18} {30}= \frac {43} {30} \)

OpenStudy (anonymous):

So is this the way I always do problems like this? What does that 64 in yellow by your name mean?

OpenStudy (swissgirl):

i am at level 64

OpenStudy (anonymous):

ok! now I have anoother question... do I need to post my next qquestion on the board, or can I just ask you? I really appreciate you helping me. This has been so hard for me. I know how to get the answere, but don't know all or how show the steps

OpenStudy (swissgirl):

u can ask it here but usually u wld close the question and then post another question

OpenStudy (anonymous):

How to find the LCM of 32 and 28

OpenStudy (swissgirl):

hmmmm shldnt it be GCF

OpenStudy (anonymous):

Is that what you want me to do?

OpenStudy (anonymous):

it says Find the LCM of 32 and 28?

OpenStudy (swissgirl):

okkkkk we will do that then

OpenStudy (swissgirl):

ok so basically we gotta list the multiples of 32 and 24

OpenStudy (anonymous):

I also have one for GCF of 72 and 78

OpenStudy (swissgirl):

So the multiples of 28 are 1*28, 2*28, 3*28, 4*28, 5*28 Multiples of 28= 28, 56,84,112,140,168,196,224 Ok now u find the mutiples of 32

OpenStudy (anonymous):

you u want me to do this

OpenStudy (anonymous):

1x32,2x32,3x32,4x32,5x32 32=32,64,66,128,160

OpenStudy (anonymous):

ok

OpenStudy (swissgirl):

ok now continue a lil more 32*6, 32*7

OpenStudy (anonymous):

6x32,7x32

OpenStudy (anonymous):

160,192

OpenStudy (swissgirl):

7*32=224

OpenStudy (anonymous):

sorry, I forgot to to send that last one

OpenStudy (swissgirl):

28= 28, 56,84,112,140,168,196,224 32=32,64,66,128,160,192,224 So what is the LCM

OpenStudy (swissgirl):

What is the lowest number that is the same?

OpenStudy (anonymous):

LCM measns least common factor, Right? So its 224

OpenStudy (swissgirl):

LCM=Lowest Common Multiple and it is 224 correct

OpenStudy (anonymous):

Great... so is this the same way I do the GCF for 72 and 78

OpenStudy (swissgirl):

Nope

OpenStudy (swissgirl):

GCF= Greatest Common Factors

OpenStudy (swissgirl):

Factors are 2 numbers that when multiplied will equal that number ok ill show u with 72

OpenStudy (anonymous):

wowwwwwwww! Really!! So do you have time to explain explain that one to me?

OpenStudy (anonymous):

OK! Thanks

OpenStudy (swissgirl):

72=1,72,2,36,3,24,4,18,6,12, 8,9 71=1,2,3,4,68,9,12,18,24,36,72

OpenStudy (swissgirl):

72=1,72,2,36,3,24,4,18,6,12, 8,9 72=1,2,3,4,68,9,12,18,24,36,72

OpenStudy (anonymous):

hold up I'm confused

OpenStudy (swissgirl):

ok well let me explain the concept with a smaller number so its easier to understand Lets take the number 6 1*6=6 2*3=6 So the factors of 6 are=1,2,3,6

OpenStudy (anonymous):

ok, let me expalin it back to you

OpenStudy (anonymous):

72=1x72=72,2x32=72,3x24=72,....etc

OpenStudy (swissgirl):

YAAAAAAAAAAAAAAAAAAAA

OpenStudy (anonymous):

wowwwwwwww

OpenStudy (anonymous):

I was a little confused

OpenStudy (swissgirl):

ok so try 78 find the factors of 78

OpenStudy (anonymous):

0k! I will try

OpenStudy (swissgirl):

well 1*78=78

OpenStudy (anonymous):

1,78,2,39,3,26

OpenStudy (anonymous):

when I got to 4 I had a remainder, so what do you do twen that happens... I stop

OpenStudy (swissgirl):

then its not a facto go onto the next one

OpenStudy (anonymous):

ummmmm... 6,13,

OpenStudy (swissgirl):

yes

OpenStudy (swissgirl):

and that is it correct

OpenStudy (swissgirl):

78=1,2,3,6,1326,39.78

OpenStudy (anonymous):

how did you get 39

OpenStudy (swissgirl):

72=1,2,3,4,68,9,12,18,24,36,72 78=1,2,3,6,13,26,39.78

OpenStudy (swissgirl):

I just copied what u gave me above

OpenStudy (swissgirl):

and u wrote 39 1,78,2,39,3,26

OpenStudy (anonymous):

ok.. so did I do it right

OpenStudy (swissgirl):

yessssssssssssss

OpenStudy (swissgirl):

72=1,2,3,4,68,9,12,18,24,36,72 78=1,2,3,6,13,26,39.78 Now we gotta find the greatest common factor What is the greatest number that is in both sets basically

OpenStudy (anonymous):

This is a little different from the other.. lol

OpenStudy (swissgirl):

it is lol. Its easy to confuse both of them

OpenStudy (anonymous):

36 and 39

OpenStudy (anonymous):

As long as I have your notes, I will be able to do this

OpenStudy (swissgirl):

nope that isnt the same number

OpenStudy (swissgirl):

we need the exact number

OpenStudy (swissgirl):

72=1,2,3,4,6,8,9,12,18,24,36,72 78=1,2,3,6,13,26,39.78

OpenStudy (anonymous):

oh .. sorry 6

OpenStudy (swissgirl):

yaaaaaaaaaaaaaaaaaaaaaaaaaa

OpenStudy (swissgirl):

There we goooooo

OpenStudy (anonymous):

Had to think about what you had really asked me

OpenStudy (anonymous):

wowwwww

OpenStudy (anonymous):

Can I save this information you have taught me?

OpenStudy (swissgirl):

hmmmmmm well if u look ontop on ur righthand corner There should be ur username, Dawson4070 in Blue. Click on it. And then a new page will open and on ur right hand side u will see something that says questions asked click on that and u will see ur question there. Try it

OpenStudy (anonymous):

yes I I see it.. but don't see anything that says for me to save it

OpenStudy (swissgirl):

No its automatically saved and if u want to find ur questions that u asked u go there

OpenStudy (swissgirl):

cuz now once ur finished u press close question and the way to get back to it is to go there

OpenStudy (anonymous):

Oh really

OpenStudy (anonymous):

will I always be avble to get you, when I come back on?

OpenStudy (swissgirl):

Ummm Not always but U can always send me a msg to see if I am on or tag me like Using the @ symbol @Dawson4070

OpenStudy (swissgirl):

But there are other ppl on here that are really smart so dont worry abt it

OpenStudy (swissgirl):

There is always someone on here to help u

OpenStudy (anonymous):

ok.. .. let me see if I can do that

OpenStudy (anonymous):

Will your messages and name be name still be in my in box once I close out out session?

OpenStudy (swissgirl):

ummm it will be but in order to find them u will have to click on the letter in the upper hand left hand side and at the bottim it says view all messages and everything will come up

OpenStudy (swissgirl):

or u can fan me and like u will be able to see if I am on or off